# Perfect pulsed inline twin-beam squeezers

### v1

@@ V.4 Double pass with arbitrary poling and polarization swap
$\bm{S}$ $=\bm{B}\begin{pmatrix}\bm{\mathcal{A}}^{T}\bm{\mathcal{A}}&0_{2N\times 2N}\\ 0_{2N\times 2N}&\bm{\mathcal{A}}^{-1}\bm{\mathcal{A}}^{-T}\\ \end{pmatrix}\bm{B}^{T}$ (210)
@@
We see that the double pass structure of Fig. 2(c) leads to input and output Schmidt modes that are equivalent and thus represents a perfect inline squeezer. All of the results above are for the SGVM regime, we now want to see if anything can be said for arbitrary signal and idler velocities.
@@ V.5 Arbitrary signal and idler group velocity
with matrix elements described above (Eqs. 42-44). The orthogonal matrix, $\bm{B}$ (Eq. 80), does not transform the matrix $\tilde{\bm{Q}}$ into block-diagonal form. One needs to consider a different transformation of the form
@@ VI Conclusion
In the SGVM regime, we find that single pass structures with and without poling lead to distinct input and output temporal modes. Furthermore, we find that these temporal modes have their real and imaginary parts swapped. By considering a frequency symmetric pump, we also find that the real and imaginary parts are flips of each other, which in the proper gauge explains the numerical results of Fig. 3
On the other hand, we find that the double pass configuration leads to equivalent input and output modes. This is due to the polarization swap which causes the input(output) modes of the second pass to be the output(input) modes of the first pass. Combining both in turn leads to an overall squeezer with the same input and output modes. Hence, not only is the double pass configuration practical for quantum computation Houde and Quesada (2023), it also serves as a perfect inline squeezer.
@@ Acknowledgements
The authors acknowledge support from the Ministère de l’Économie et de l’Innovation du Quèbec and the Natural Sciences and Engineering Research Council of Canada. The authors thank W. McCutcheon for insightful discussions concerning Bloch-Messiah decomposition and its implementation in Python, J.E. Sipe for insightful discussions, and Karthik Chinni for providing feedback on the manuscript.

### v2

@@ V.4 Double pass with arbitrary poling and polarization swap
$\bm{S}$ $=\bm{B}\begin{pmatrix}\bm{\mathcal{A}}^{T}\bm{\mathcal{A}}&0_{2N\times 2N}\\ 0_{2N\times 2N}&\bm{\mathcal{A}}^{-1}\bm{\mathcal{A}}^{-T}\\ \end{pmatrix}\bm{B}^{T}$
@@
We see that the double pass structure of Fig. 2(c) leads to input and output Schmidt modes that are equivalent and thus represents a perfect inline squeezer. This conclusion only holds in the ideal case where both passes are identical. We now want to numerically study how robust the double pass structure is to variations in gain between the two passes.
V.4.1 Robustness of double pass structure
In a typical experimental setting, several things could occur in between the passes which would effectively lower the gain level in the second pass. Such effects could be caused by losses or by some phase buildup between the modes which causes the phase-matching condition to no longer be fully met. One could also imagine a double pass structure where each pass is a unique nonlinear region which could have variations in their nonlinear strengths due to the fabrication process, allowing for a possibly higher or lower gain in the second pass.
To study these effects, we consider a setup where we fix the level of gain in the first pass and allow the second pass to vary. We choose the parameters in such a way that when both passes give rise to the same gain, the average number of signal photons after the double pass is $\langle N_{S}\rangle=5$. We then vary the gain in the second pass such that the average number of signal photons after the double pass varies from $\langle N_{S}\rangle=2.5$ to $\langle N_{S}\rangle=7.5$ which is a difference in expected gain of ±$\pm 50\%$. We numerically compute the Bloch-Messiah decomposition as we sweep through the gain of the second pass and calculate the fidelity between the input Schmidt mode $\rho^{in}_{S}(\omega)$ and the output Schmidt mode $\rho^{out}_{S}(\omega)$ which we define as
[…]
Figure 4 shows the fidelity as a function of the overall gain as characterized by $\langle N_{S}\rangle$. We see that even if the expected gain varies by ±$\pm 50\%$, the fidelity remains quite high ($>0.99$). This shows that the double pass structure is a robust perfect inline squeezer. All of the results above are for the SGVM regime, we now want to see if anything can be said for arbitrary signal and idler velocities.
Figure 4: Fidelity between the input and output Schmidt modes for a double pass structure where the gain in the second pass is allowed to vary. We choose parameters such that when both passes have the same gain the average number of signal photons is $\langle N_{S}\rangle=5$. We allow for total variations of the average number of signal photons on the order of ±$\pm 2.5$. We see that the fidelity between input and output Schmidt modes remains very high.
@@ V.5 Arbitrary signal and idler group velocity
with matrix elements described above (Eqs. 30-32). The orthogonal matrix, $\bm{B}$ (Eq. 40), does not transform the matrix $\tilde{\bm{Q}}$ into block-diagonal form. One needs to consider a different transformation of the form
@@ VI Conclusion
In the SGVM regime, we find that single pass structures with and without poling lead to distinct input and output temporal modes. Furthermore, we find that these temporal modes have their real and imaginary parts swapped. By considering a frequency symmetric pump, we also find that the real and imaginary parts are flips of each other, which in the proper gauge explains the numerical results of Fig. 3. Experimentally, one could imagine using a phase-shifter to correct the phases before sending the signal back into the single pass structure. However, for the pump pulse time-scales considered (femtoseconds) no such device exists. State of the art phase-shifters work up to the picoseconds regime. Furthermore, the single pass Schmidt mode structure is highly gain dependentHoude and Quesada (2023) and it would require extensive calibration to ensure that the phase-shifter is properly correcting for the phase mismatch. Even if one could use a phase-shifter it would highly convolute the experimental design due to this calibration. Phase-shifters also introduce significant noise in the system which would severely affect the squeezing properties of the setup. As such, the single pass configuration does not serve as a perfect inline squeezer.
On the other hand, we find that the double pass configuration leads to equivalent input and output modes. This is due to the polarization swap which causes the input(output) modes of the second pass to be the output(input) modes of the first pass. Combining both in turn leads to an overall squeezer with the same input and output modes. Hence, not only is the double pass configuration practical for quantum computation Houde and Quesada (2023), it also serves as a perfect inline squeezer. In section V.4.1, we numerically study how relaxing the assumption that both passes have the same level of gain affects the input and output temporal modes and find that the double pass structure is a robust perfect inline squeezer.
@@ Acknowledgements
The authors acknowledge support from the Ministère de l’Économie et de l’Innovation du Québec and the Natural Sciences and Engineering Research Council of Canada. This work has been funded by the European Union’s Horizon Europe Research and Innovation Programme under agreement 101070700 project MIRAQLS. The authors thank W. McCutcheon for insightful discussions concerning Bloch-Messiah decomposition and its implementation in Python, J.E. Sipe and Guillaume Thekkadath for insightful discussions, and Karthik Chinni for providing feedback on the manuscript.